# On the Lipschitz Continuity of Set Aggregation Functions and Neural Networks for Sets

### v3

@@ D.3 Stability under Perturbations of Input Multisets
Each document of the Polarity dataset is represented as a multiset of word vectors. The word vectors are obtained from a publicly available pre-trained model (Mikolov et al., 2013). We randomly split the dataset into training, validation, and test sets with a $80:10:10$ split ratio. The $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ models consist of an MLP which transforms the representations of the words, an aggregation function (mean and max, respectively) and a second MLP which produces the output. Both MLPs consist of two hidden layers. The ReLU function is applied to the outputs of the first layer. The hidden dimension size is set to $64$ for all hidden layers. The model is trained for $20$ epochs by minimizing the cross-entropy loss function with the Adam optimizer and a learning rate of $0.001$. At the end of each epoch, we compute the performance of the model on the validation set, and we choose as our final model the one that achieved the smallest loss on the validation set.
@@ E.1 Lipschitz Constant of Aggregation Functions
We next provide some additional empirical results that validate the findings of Theorem 3.1 (i. e., Lipschitz constants of aggregation functions). We experiment with the Polarity dataset. Figure 4 visualizes the relationship between the output of the three considered distance functions and the Euclidean distance of the aggregated representations of multisets of documents from the Polarity dataset. Since the Polarity dataset consists of documents which might differ from each other in the number of terms, by Theorem 3.1 we can derive upper bounds only for $3$ out of the $9$ combinations of distance functions for multisets and aggregation functions. As expected, the Lipschitz bounds (dash lines) upper bound the Euclidean distance of the outputs of the aggregation functions. With regards to the tightness of the bounds, we observe that the bounds that are associated with the mean and sum functions are tighter than the one associated with the max function. The correlations between the distances of multisets and the Euclidean distances of the aggregated representations of multisets are relatively low in most of the cases. All three distance functions are mostly correlated with the aggregation functions with which they are related via Theorem 3.1. The highest correlation is achieved between EMD and the mean function ($r=0.89$).
@@ E.2 Upper Bounds of Lipschitz Constants of Neural Networks for Sets
We next provide some additional empirical results that validate the findings of Theorem 3.4 (i. e., upper bounds of Lipschitz constants of neural networks for sets). The results are shown in Figure 5. Since the Polarity dataset consists of documents which might differ from each other in the number of terms, by Theorem 3.4 we can derive upper bounds only for $2$ out of the $9$ combinations of distance functions for multisets and aggregation functions. While the dash lines upper bound the Euclidean distance of the outputs of the aggregation functions, both bounds are relatively loose on the Polarity dataset. The correlations between the distances of the representations produced by the neural network for the multisets and the distances produced by distance functions for multisets are much lower than those of the previous experiments. The highest correlation is equal to $0.55$ (between the neural network that utilizes the mean aggregation function and EMD), while there are even negative correlations. This is not surprising since for most of the combinations, there are no upper bounds on the Lipschitz constant of the corresponding neural networks.
@@ E.4 Generalization under Distribution Shifts
To evaluate the generalization of the two Lipschitz continuous models ($\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$) under distribution shifts, we also experiment with the Amazon review dataset (Blitzer et al., 2007). The dataset consists of product reviews from Amazon for four different types of products (domains), namely books, DVDs, electronics and kitchen appliances. For each domain, there exist $2,000$ labeled reviews (positive or negative) and the classes are balanced. We construct $4$ adaptation tasks. In each task, the $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ models are trained on reviews for a single type of products and evaluated on all domains.
@@ E.5 Predictive Performance of Neural Networks for Sets
We next evaluate the $\textsc{NN}_{\textsc{sum}}$, $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ models on four classification and regression datasets, namely ModelNet40, Polarity, IMDB and IMDB-BINARY. The first two datasets are described in section 4. IMDB contains movie reviews from the IMDb database (Maas et al., 2011). The targets are the ratings that accompany the reviews ($10$ different values). We treat this task as a regression problem. IMDB-BINARY is a standard graph classification dataset (Yanardag and Vishwanathan, 2015), commonly used for evaluating graph kernels and graph neural networks. Each graph of the IMDB-BINARY dataset was represented as a multiset of the degrees of its nodes. The results are illustrated in Table 3. Each experiment was repeated $5$ times with different random seeds, and for each dataset we report average accuracy (for ModelNet40, Polarity and IMDB-BINARY) or average root mean square error (for IMDB) on the dataset’s test set and the corresponding standard deviation.
@@
Polarity consists of short reviews (average number of terms = $20$). Therefore, whether a review is positive or negative depends primarily on the presence of one or a few terms that indicate sentiment. These terms can be considered extreme elements, and the Hausdorff distance relies on such extreme elements when comparing its inputs. This distance function thus seems to be suitable for this task and potentially explains why $\textsc{NN}_{\textsc{max}}$ is the best-performing model on this dataset.

### v4

@@ D.3 Stability under Perturbations of Input Multisets
Each document of the Polarity dataset is represented as a multiset of word vectors. The word vectors are obtained from a publicly available pre-trained model (Mikolov et al., 2013). We randomly split the dataset into training, validation, and test sets with a $80:10:10$ split ratio. The $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ models consist of an MLP which transforms the representations of the words, an aggregation function (mean and max, respectively) and a second MLP which produces the output. Both MLPs consist of two hidden layers. The ReLU function is applied to the outputs of the first layer. The hidden dimension size is set to $64$ for all hidden layers. The model is trained for 2020 epochs by minimizing the cross-entropy loss function with the Adam optimizer and a learning rate of $0.001$. At the end of each epoch, we compute the performance of the model on the validation set, and we choose as our final model the one that achieved the smallest loss on the validation set.
@@ E.1 Lipschitz Constant of Aggregation Functions
We next provide some additional empirical results that validate the findings of Theorem 3.1 (i. e., Lipschitz constants of aggregation functions). We experiment with the Polarity dataset. Figure 4 visualizes the relationship between the output of the three considered distance functions and the Euclidean distance of the aggregated representations of multisets of documents from the Polarity dataset. Since the Polarity dataset consists of documents which might differ from each other in the number of terms, by Theorem 3.1 we can derive upper bounds only for 33 out of the 99 combinations of distance functions for multisets and aggregation functions. As expected, the Lipschitz bounds (dash lines) upper bound the Euclidean distance of the outputs of the aggregation functions. With regards to the tightness of the bounds, we observe that the bounds that are associated with the mean and sum functions are tighter than the one associated with the max function. The correlations between the distances of multisets and the Euclidean distances of the aggregated representations of multisets are relatively low in most of the cases. All three distance functions are mostly correlated with the aggregation functions with which they are related via Theorem 3.1. The highest correlation is achieved between EMD and the mean function ($r=0.89$).
@@ E.2 Upper Bounds of Lipschitz Constants of Neural Networks for Sets
We next provide some additional empirical results that validate the findings of Theorem 3.4 (i. e., upper bounds of Lipschitz constants of neural networks for sets). The results are shown in Figure 5. Since the Polarity dataset consists of documents which might differ from each other in the number of terms, by Theorem 3.4 we can derive upper bounds only for 22 out of the 99 combinations of distance functions for multisets and aggregation functions. While the dash lines upper bound the Euclidean distance of the outputs of the aggregation functions, both bounds are relatively loose on the Polarity dataset. The correlations between the distances of the representations produced by the neural network for the multisets and the distances produced by distance functions for multisets are much lower than those of the previous experiments. The highest correlation is equal to $0.55$ (between the neural network that utilizes the mean aggregation function and EMD), while there are even negative correlations. This is not surprising since for most of the combinations, there are no upper bounds on the Lipschitz constant of the corresponding neural networks.
@@ E.4 Generalization under Distribution Shifts
To evaluate the generalization of the two Lipschitz continuous models ($\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$) under distribution shifts, we also experiment with the Amazon review dataset (Blitzer et al., 2007). The dataset consists of product reviews from Amazon for four different types of products (domains), namely books, DVDs, electronics and kitchen appliances. For each domain, there exist $2,000$ labeled reviews (positive or negative) and the classes are balanced. We construct 44 adaptation tasks. In each task, the $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ models are trained on reviews for a single type of products and evaluated on all domains.
@@ E.5 Predictive Performance of Neural Networks for Sets
We next evaluate the $\textsc{NN}_{\textsc{sum}}$, $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ models on four classification and regression datasets, namely ModelNet40, Polarity, IMDB and IMDB-BINARY. The first two datasets are described in section 4. IMDB contains movie reviews from the IMDb database (Maas et al., 2011). The targets are the ratings that accompany the reviews ($10$ different values). We treat this task as a regression problem. IMDB-BINARY is a standard graph classification dataset (Yanardag and Vishwanathan, 2015), commonly used for evaluating graph kernels and graph neural networks. Each graph of the IMDB-BINARY dataset was represented as a multiset of the degrees of its nodes. The results are illustrated in Table 3. Each experiment was repeated 55 times with different random seeds, and for each dataset we report average accuracy (for ModelNet40, Polarity and IMDB-BINARY) or average root mean square error (for IMDB) on the dataset’s test set and the corresponding standard deviation.
@@
Polarity consists of short reviews (average number of terms = 2020). Therefore, whether a review is positive or negative depends primarily on the presence of one or a few terms that indicate sentiment. These terms can be considered extreme elements, and the Hausdorff distance relies on such extreme elements when comparing its inputs. This distance function thus seems to be suitable for this task and potentially explains why $\textsc{NN}_{\textsc{max}}$ is the best-performing model on this dataset.